\begin{thebibliography}{99}

\bibitem{Korteweg 95 a}
Korteweg DJ \& de Vries G (1895) On the change of form of long waves advancing in a rectangular canal and on a new type of long stationary waves {\it Philosophical Magazine, 5th series} {\bf 39} 422--443

\bibitem{Hirota 71 a}
Hirota R (1971) Exact solution of the Korteweg--de Vries equation for multiple collisions of solitons {\it Physical Review Letters} {\bf 27} 1192--1194

\bibitem{Hirota 73 a}
Hirota R (1973) Exact envelope-soliton solutions of a nonlinear wave equation {\it Journal of Mathematical Physics} {\bf 14} 805--809

\bibitem{Hirota 73 b}
Hirota R (1973) Exact N-soliton solutions of the wave equation of long waves in shallow-water and in nonlinear lattices {\it Journal of Mathematical Physics} {\bf 14} 810--814

\bibitem{Hirota 76 a}
Hirota R \&  Satsuma J (1976) N-soliton solution of the K-dV equation with loss and nonuniformity terms {\it Journal of the Physical Society of Japan} {\bf 41} 2141--2142

\bibitem{Hirota 74 a}
Hirota R (1974) A new form of B\"acklund transformation and its relation to the inverse scattering problem {\it Progress in Theoretical Physics} {\bf 52}  1498--1512

\bibitem{Gui-2013}
Gui G, Liu Y, Olver PJ \& Qu C (2013) Wave-breaking and peakons for a modified Camassa-Holm equation {\it Communications in Mathematical Physics} {\bf 319} 731--759

\bibitem{Fuch-1996}
    Fuchssteiner B (1996) Some tricks from the symmetry-toolbox for nonlinear equations: generalizations of the Camassa-Holm equation
    {\it Physica D} {\bf 95}  229--243

\bibitem{Fokas1995}
Fokas AS (1995) The Korteweg-de Vries equation and beyond {\it Acta Applicandae Mathematicae}  {\bf 39} 295--305

\bibitem{Xia-2018}
Xia B, Qiao Z \& Li JB (2018) An integrable system with peakon, complex peakon, weak kink, and kink-peakon interactional solutions
{\it Communications in Nonlinear Science and Numerical Simulation} {\bf 63}  292--306

\bibitem{Fu-2013}
Fu Y, Gui G, Qu C \& Liu Y (2013) On the Cauchy problem for the integrable Camassa-Holm type equation with cubic nonlinearity {\it Journal of Differential Equations} {\bf 255} 1905--1938

\bibitem{Lie74a}
Lie S (1874) \"Uber die Integration durch bestimmte Integrale von einer Klasse linear partieller Differentialgleichungen {\it Mathemathischen Annallen} {\bf 8} 328-368

\bibitem{Olver-86}
Olver PJ (1986) {\it Applications of Lie Groups to Differential Equations} (Springer, New York)

\bibitem{Leach 88 a}
Leach PGL, Feix MR \& Bouquet S (1988) Analysis and solution of a nonlinear second-order differential equation through rescaling and through a dynamical point of view {\it Journal of Mathematical Physics} {\bf 29} 2563--2569

\bibitem{Leach 99}
Leach PGL, Govinder KS \& Abraham-Shrauner B (1999) Symmetries of first integrals and their associated differential equations {\it Journal of Mathematical Analysis and Application} {\bf 235} 58--83

\bibitem{Leach 12 a}
Leach PGL, Govinder KS \& Andriopoulos K (2012) Hidden and not so hidden symmetries {\it  Journal of Applied Mathematics} {\bf 2012} Article ID 890171, (doi:10.1155/2011/890171)

\bibitem{Andriopoulos09a}
Andriopoulos K, Dimas S, Leach PGL \& Tsoubelis D (2009) On the systematic approach to the classification of differential equations by group theoretical methods {\it Journal of Computational and Applied Mathematics} {\bf 230}  224--232 (DOI: 10.1016/j.cam.2008.11.002)

\bibitem{Tamizhmani 14 a}
 Tamizhmani KM, Krishnakumar K \& Leach PGL (2014) Algebraic resolution of equations of the Black-Scholes type with arbitrary time-dependent parameters {\it Applied Mathematics and Computation} {\bf 247} 115--124

\bibitem{Krishnakumar 14 b}
Tamizhmani KM, Sinuvasan R, Krishnakumar K \& Leach PGL (2014) Some symmetry properties of the Riccati Differential Sequence and its integrals  {\it Afrika Matematika} 1--6


\bibitem{Winternitz09a}
Rodr\'{\i}guez MA, Tempesta P \& Winternitz P (2009) Symmetry reduction and superintegrable Hamiltonian systems. {\it Journal of Physics: Conference Series} {\bf 175}.

\bibitem{Ablowitz 80a}
Ablowitz MJ, Ramani A \& Segur H (1980) A connection between nonlinear evolution equations and ordinary differential equations of P-type I {\it Journal of Mathematical Physics} {\bf 21} 715--721

\bibitem{Ablowitz 80b}
Ablowitz MJ, Ramani A \& Segur H (1980) A connection between nonlinear evolution equations and ordinary differential equations of P-type II {\it Journal of Mathematical Physics} {\bf 21} 1006--1015

\bibitem{Ramani 89}
Ramani A, Grammaticos B \& Bountis T (1989) The Painlev\'{e} Property and singularity analysis of integrable and non-integrable systems {\it Physics Reports} {\bf 108} 159--245

\bibitem{Feix 97 a}
Feix MR, Geronimi C, Cair\'{o} L, Leach PGL, Lemmer RL \& Bouquet S\'E (1997) Right and left Painlev\'{e} series for ordinary differential equations invariant under time translation and rescaling {\it Journal of Physics A: Mathematical and General} {\bf 30} 7437--7461

\bibitem{Andriopoulos 06 a}
Andriopoulos K \& Leach PGL (2006) An interpretation of the presence of both positive and negative nongeneric resonances in the singularity analysis {\it Physics Letters A} {\bf 359} 199--203

\bibitem{Lemmer 93 a}
Lemmer RL \& Leach PGL (1993) The Painlev\'{e} test, hidden symmetries and the equation $y^{\prime\prime}+ yy^{\prime}+ky^{3} = 0$ {\it Journal of Physics A: Mathematical and General} {\bf 26}  5017--5024

\bibitem{G. Adomian88}
Adomian G (1988) A review of the decomposition method in applied mathematics {\it Journal of Mathematical Analysis and Applications} {\bf 135} 501-544

\bibitem{Wazwaz}
Wazwaz AM (2006) Solitary wave solutions for modified forms of Degasperis-Procesi and Camassa-Holm Equations. {\it Physics Letter A} {\bf 352} 500--504

\bibitem{Majeed}
Majeed A. Yousif, Bewar A. Mahmood \& Fadhil H. Easif (2015) A New Analytical Study of Modified  Camassa-Holm and Degasperis-Procesi  Equations {\it American Journal of Computational Mathematics} {\bf 5} 267--273 (doi:10.1063/1.1514385)

\bibitem{Winternitz02}
Gravel S \& Winternitz P (2002) Superintegrability with third-order integrals in quantum and classical mechanics. {\it Journal of Mathematical Physics} {\bf 43} 5902-5912 (doi:10.1063/1.1514385)

\bibitem{Winternitz09b}
Winternitz P \& Yurduşen I (2009) Integrable and superintegrable systems with spin in three-dimensional Euclidean space. {\it Journal of Physics A: Mathematical and Theoritical} {\bf 42} 385203 (doi:10.1088/1751-8113/42/38/385203)

\bibitem{Olver-1996}
Olver PJ \& Rosenau P (1996) Tri-Hamiltonian duality between solitons and solitary-wave solutions having compact support {\it Physics Review E} {\bf 53}  1900--1906

\bibitem{Qiao-2006}
Qiao Z (2006) A new integrable equation with cuspons and W/M-shape-peaks solitons {\it Journal of Mathematical Physics} {\bf 47}  112701

\bibitem{Himonas-2014}
Himonas A \& Mantzavinos D (2014) The Cauchy problem for the Fokas-Olver-Rosenau-Qiao equation {\it Nonlinear Analysis} {\bf 95}  499--529

\bibitem{Yang-2017}
Yang M, Li Y \& Zhao Y (2017) On the Cauchy problem of generalized Fokas-Olver-Resenau-Qiao equation {\it Applicable Analysis} http://dx.doi.org/10.1080/00036811.2017.1359565.

\bibitem{Qiao-2011}
Qiao Z \& Li X (2011) An integrable equation with nonsmooth solitons {\it Theoretical and Mathematical Physics} {\bf 267}  584--589

\bibitem{Camassaholm}
Camassa R \& Holm D (1993) An integrable shallow water equation with peaked solitons {\it Physics Review Letter} {\bf 71} 1661--1664

\bibitem{ConstantinLannes}
Constantin A \& Lannes D (2009) The hydrodynamical relevance of the Camassa-Holm and Degasperis-Procesi equations {\it Archive for Rational Mechanics and Analysis} {\bf 192} 165--186

\bibitem{FuchssteinerFokas}
Fuchssteiner B \& Fokas A (1981/1982) Symplectic structures, their Bäcklund transformations and hereditary symmetries {\it Physica D} {\bf 4} 47--66

\bibitem{ConstantinStrauss}
Constantin A \& Strauss WA (2000) Stability of a class of solitary waves in compressible elastic rods {\it Physics Letter A} {\bf 270} 140--148

\bibitem{Dimas05a}
Dimas S \& Tsoubelis D (2005) SYM: A new symmetry-finding package for Mathematica {\it Group Analysis of Differential Equations} Ibragimov NH, Sophocleous C \& Damianou PA edd (University of Cyprus, Nicosia) 64-70  See also  http : // www.math.upatras.gr/~spawn

\bibitem{Dimas06a}
Dimas S \& Tsoubelis D (2006) A new Mathematica-based program for solving overdetermined systems of PDEs {\it 8th International Mathematica Symposium} (Avignon, France)

\bibitem{Dimas08a}
Dimas S  (2008) {\it Partial Differential Equations, Algebraic Computing and Nonlinear Systems} (Thesis: University of Patras, Patras, Greece)

\bibitem{GaoLiuQu}
Zihua Gao,Xiaochuan Liu, Xingxing Liu \& Changzheng Qu (2019) Stability of Peakons for the generalized Modified Camassa-Holm equation {\it Journal of Differential equations} {\bf 266} 7749--7779

\bibitem{Painleve97}
Painlev\'e P  (1973) Le\c{c}ons sur la th\'eorie analytique des \'equations diff\'erentielles. (Le\c{c}ons de Stockholm, 1895) (Hermann, Paris, 1897). Reprinted, {O$\!$euvres de Paul Painlev\'e}, vol.~I, \'Editions du CNRS, Paris.

\bibitem {Painleve00}
Painlev\'e P (1900) M\'emoire sur les \'equations diff\'erentielles du second ordre dont l'int\'egrale g\'en\'erale est uniforme. {\it Bulletin of the Mathematical Society of France} {\bf 28} 201-265

\bibitem {Painleve02}
Painlev\'e P (1902) Sur les \'equations diff\'erentielles du second ordre et d'ordre sup\'erieur dont l'int\'egrale g\'en\'erale est uniforme {\it Acta Mathematica} {\bf 25} 1-85

\bibitem{Krishnakumar 16}
Krishnakumar K (2016) {\it A study of symmetries, reductions and solutions of certain classes of differential equations} (Thesis: Pondicherry Central University, Puducherry, India)

\end{thebibliography}

\documentclass[a4paper, 12pt]{article}
\usepackage{amsmath, amssymb, graphics}
\newcommand{\mathsym}[1]{{}}
\newcommand{\unicode}[1]{{}}

\catcode`\@=11

\@addtoreset{equation}{section} 

\def\dddot#1{\mathinner{\buildrel\vbox{\kern5pt\hbox{...}}\over{#1}}}
\def\ddddot#1{\mathinner{\buildrel\vbox{\kern5pt\hbox{....}}\over{#1}}}

\begin{document}

\begin {center}
{\Large Symmetries and Integrability of Modified Camassa-Holm Equation with an Arbitrary Parameter}\\[3 mm]
{\small A Durga Devi $ ^ 1$, K Krishnakumar$ ^ {2*}$ , R Sinuvasan$ ^ 3$ and PGL Leach$ ^ 4$\\[3 mm]
$^ 1 $ Department of Physics, Srinivasa Ramanujan Centre, SASTRA Deemed to be University, Kumbakonam 612 001, India.\\$^ 2 $ Department of Mathematics, Srinivasa Ramanujan Centre, SASTRA Deemed to be University, Kumbakonam 612 001, India.\\$^ 3 $ Department of Mathematics, SASTRA Deemed to be University, Thanjavur 613 401, India.\\$^4 $Department of Mathematics, Durban University of Technology, PO Box 1334,\\ Durban ~4000,~ Republic of South Africa.}
\end{center}
\begin {abstract}
We study the symmetry and integrability of a modified Camassa-Holm Equation (MCH), with an arbitrary parameter $k,$ of the form $$u_{t}+k(u-u_{xx})^2u_{x}-u_{xxt}+(u^{2}-{u_{x}}^2)(u_{x}-u_{xxx})=0.$$ By using Lie point symmetries we reduce the order of the above equation and also we obtain interesting novel solutions for the reduced ordinary differential equations. Finally we apply the Painlev\'e Test to the resultant nonlinear ordinary differential equation.
\end {abstract}
{\bf MSC Numbers:}  17B80; 34M15; 58J70\\
{\bf PACS Numbers:} 02.20.Sv; 02.30.Hq; 02.30.Ik\\
{\bf Keywords:} Modified Camassa-Holm Equation (MCH), Symmetry, Painlev\'e Test, Integrability.

\section {Introduction}
The linear and nonlinear partial differential equations have encompassed the landscape of many disciplines such as chemical, physical, biological sciences, engineering and mathematics due to its potential ability to provide adequate understanding of the underlying systems.
There are many wave equations in engineering and physics that are quite helpful and relevant such as the Korteweg–de Vries (KdV) equation, nonlinear coupled KdV equations, Modified Camassa-Holm (MCH) equation and so on \cite{Korteweg 95 a, Hirota 71 a, Hirota 73 a, Hirota 73 b, Hirota 76 a, Hirota 74 a, Gui-2013, Fuch-1996, Fokas1995, Xia-2018, Fu-2013}. In general, the closed- form solutions of the differential equations are very hard to achieve in most instances.

Numerous systematic techniques still have been developed to find the solutions of differential equations in recent years. Especially, Lie Symmetry Analysis \cite{Lie74a, Olver-86, Leach 88 a, Leach 99, Leach 12 a, Andriopoulos09a, Tamizhmani 14 a, Krishnakumar 14 b, Winternitz09a}, Painlev\'e Analysis \cite{Ablowitz 80a, Ablowitz 80b, Ramani 89, Feix 97 a, Andriopoulos 06 a, Lemmer 93 a}, Adomian decomposition method \cite{G. Adomian88, Wazwaz}, the variational iteration method, the homotopy analysis method,  the homotopy perturbation method and variational homotopy perturbation method have been utilized successfully to solve and study the nature of integrability of partial  differential equations \cite{Majeed}. Among these methods the symmetry analysis and Painlev\'e Analysis have been used by researchers effectively \cite{Winternitz09a, Winternitz02, Winternitz09b}.

\strut\hfill

Recently researchers have been inspired to study the integrability and solutions of various Camassa-Holm (CH) types of equations with cubic or higher-order nonlinearities. Olver and Rosenau \cite{Olver-1996} and Fuchssteiner \cite{Fuch-1996} derived the well known modified Camassa-Holm (MCH) equation:
\begin{eqnarray}
y_{t}+u y_{x}+2u_{x}y=0,\nonumber\\
y=u-u_{xx}
\label{mch}
\end{eqnarray}
 by employing the triHamiltonian duality approach to the biHamiltonian representation of the MKDV equation. The functions $u$ and $y$ represent, the velocity of the fluid and its potential density \cite{Qiao-2006}. The MCH equation is a water-wave equation and is a suitable approximation of the incompressible irrotational Euler system. Consequently Qiao \cite{Qiao-2006} discussed the nature of integrability and the structure of solutions of the Modified Camassa Holm equation.  In some literature the modified Camassa Holm equation is otherwise state as FORQ equation \cite{Himonas-2014, Yang-2017}.

 \strut\hfill

	The MCH equation is completely Integrable \cite{Olver-1996}.  It has a biHamiltonian structure and also admits a Lax pair \cite{Qiao-2011} and hence may be solved by the inverse scattering transform method. The MCH equation can be viewed as a cubic extension of the well-known Camassa-Holm (CH) equation which was proposed as a model describing the unidirectional propagation of shallow
water waves \cite{Camassaholm, ConstantinLannes, FuchssteinerFokas} and axially symmetric waves in hyperelastic rods \cite{ConstantinStrauss}.

\strut\hfill

	The purpose of this work is to study the symmetry and integrability of MCH equation with an arbitrary parameter which is given by
 \begin{eqnarray}
u_{t}+k(u-u_{xx})^2u_{x}-u_{xxt}+(u^{2}-{u_{x}}^2)(u_{x}-u_{xxx})=0,
\label{mcha2}
\end{eqnarray}
 where parameter $k \in R$ characterizes the magnitude of the linear dispersion.

 \strut\hfill

 We split this article into two sections. In the first section we discuss Lie point symmetries and reductions of order for (\ref{mcha2}). Lie symmetry is an important tool to provide a systematic method for finding the solution of the given differential equations. In which, we reduce the order of the ordinary or partial differential equations by using the symmetry. Hence, Lie symmetries plays a major role in it. Lie symmetries can be used for the classification of differential equations. Throughout this literature the Mathematica add-on Sym \cite{Dimas05a, Dimas06a, Dimas08a} is used to compute the symmetries. In the second section, we study the Painlev\'e Analysis for equation (\ref{mcha2}) and we able to prove that the equation (\ref{mcha2}) is integrable.

\strut\hfill

\section{The Symmetries of modified Camassa-Holm equation}
Qiao \cite{Qiao-2006} discussed the equation
\begin{eqnarray}
u_{t}+k(u-u_{xx})^2u_{x}-u_{xxt}+(u^{2}-{u_{x}}^2)(u_{x}-u_{xxx})=0,
\label{mcha}
\end{eqnarray}
which is integrable only when $k=2$. The author also suggested to examine the integrability of the above equation for all possible values of $k$. Therefore we propose a more general form of (\ref{mcha}) by taking $k$ as $k(t)$ which is a function of $t$. Hence the modified Camassa-Holm equation (MCH) with arbitrary function, $k(t)$, \cite{Qiao-2006, GaoLiuQu} can be represented as
\begin{eqnarray}
u_{t}+k(t)(u-u_{xx})^2u_{x}-u_{xxt}+(u^{2}-{u_{x}}^2)(u_{x}-u_{xxx})=0,
\label{mchaa}
\end{eqnarray}
when $k(t)$ is an arbitrary function of $t$. Equation $(\ref{mchaa})$ has a symmetry $\partial_x$. Also we have arrived at the trivial solution $u(x,t)=C$ to the equation $(\ref{mchaa})$.
To find the nontrivial solutions, we choose $k(t)=k$ as an arbitrary parameter. Therefore the equation becomes
 \begin{eqnarray}
u_{t}+k(u-u_{xx})^2u_{x}-u_{xxt}+(u^{2}-{u_{x}}^2)(u_{x}-u_{xxx})=0.
\label{mch}
\end{eqnarray}

The symmetries of equation(\ref{mch}), when $k$ is an arbitrary parameter, are
\begin{eqnarray}
\textbf{X}_{1}&=&\partial_{x}\\
\textbf{X}_{2}&=&\partial_{t}\\
\textbf{X}_{3}&=&2t\partial_{t}-u\partial_{u}.
\end{eqnarray}
The Lie Brackets of the Lie symmetries of equation (\ref{mch}) are\\
\begin{center}
\begin{tabular}{|c|ccc|}
$\left[ \textbf{X}_{I},\textbf{X}_{J}\right] $ & $\textbf{X}_{1}$ & $\textbf{X}_{2}$ & $\textbf{X}_{3}$  \\
$\textbf{X}_{1}$ & $0$ & $0$ & $0$   \\
$\textbf{X}_{2}$ & $0$ & $0$ & $2\textbf{X}_{2}$ \\
$\textbf{X}_{3}$ & $0$ & $-2\textbf{X}_{2}$ & $0$. \\
\end{tabular}
\end{center}
Now we discuss the travelling-wave solution of equation (\ref{mch}) by taking the linear combination of $\textbf{X}_{1}$ and $\textbf{X}_{2}$.  Therefore we can represent the linear combination of these two symmetries as $\textbf{X}_{4}=\partial_{t}+C \partial_{x}$.  The corresponding canonical variables are $r=x-Ct$ and $u(x,t)=Q(r)$. Based on these canonical variables the equation($\ref{mch})$ can be reduced to
\begin{eqnarray}
 Q^{\prime\prime\prime} (C-Q^2+{Q^{\prime}}^2)+k(Q^{\prime\prime}
 -2Q)Q^{\prime}Q^{\prime\prime}-{Q^{\prime}}^3-(C-Q^2-kQ^2)Q^{\prime}=0, \label{red2}
\end{eqnarray}
where  $Q$ is the function of the new independent variable $r$.
The symmetry of the above equation(\ref{red2}) is $\partial_{r}$ for which the canonical variables are $Q(r)=Q$ and $Q^{\prime}=W(Q)$.  Based on these canonical variables the equation, (\ref{red2}), can be reduced to
\begin{eqnarray}
(C-Q^2+W^2)WW^{\prime\prime}+(C-Q^2+W^2+kW^2){W^{\prime}}^2
-2kQWW^{\prime}&&\nonumber\\-W^2+(k+1)Q^2-C=0,\label{redsecord}
\end{eqnarray}
where $W$ is a function of $Q$.  When we find the symmetries of the equation, ({\ref{redsecord}}), we arrive the following two possible cases.\\

\strut\hfill

{\bf Case:1}

 If $k\neq-2$, Then the symmetries of the equation, ({\ref{redsecord}}), are
\begin{eqnarray}
\Gamma_{1}&=&\partial_{Q}+\frac{Q}{W}\partial_{W}\\
\Gamma_{2}&=&\frac{(C-Q^{2}+W^{2})^{\frac{-k}{2}}}{W}\partial_{W}\\
\Gamma_{3}&=&\frac{Q(C-Q^{2}+W^{2})^{\frac{-k}{2}}}{W}\partial_{W}\\
\Gamma_{4}&=&\frac{(C-Q^{2}+W^{2})}{(k+2)W}\partial_{W}\\
\Gamma_{5}&=&2{Q}\partial_{Q}+\frac{(C+(3+2k)Q^2+W^2)}{(k+2)W}\partial_{W}\\
\Gamma_{6}&=&{Q}^{2}\partial_{Q}+\frac{Q(C+(k+1)Q^2+W^2)}{(k+2)W}\partial_{W}\\
\Gamma_{7}&=&\frac{(C-Q^2+W^{2})^{\frac{k+2}{2}}}{k+2}\partial_{Q}+\frac{Q(C-Q^2+W^{2})^{\frac{k+2}{2}}}{(k+2)W}\partial_{W}\\
\Gamma_{8}&=&\frac{Q(C-Q^2+W^{2})^{\frac{k+2}{2}}}{k+2}\partial_{Q}+
\frac{(C-Q^2+W^{2})^{\frac{k+2}{2}}(C+(k+1)Q^2+W^2)}{(k+2)^2W}{\partial_{W}}.\nonumber\\
\end{eqnarray}

When we use $\Gamma_{1}$, the solution of equation, ({\ref{redsecord}}), is given by
\begin{equation}
W=\pm \sqrt{Q^2+2I_{1}},
\end{equation}
where $I_{1}$ is the constant of integration.
From $\Gamma_{4}$, the solution is
\begin{equation}
W=\pm \sqrt{(Q^2-C)}.
\end{equation}

By using $\Gamma_{5}$, the solution is
\begin{equation}
W=\pm\sqrt{Q^2-C+(2Q(k+2))^{\frac{1}{k+2}}I_{2}},
\end{equation}
where $I_2$ is the constant of integration, with $(2Q(k+2))^{\frac{2}{k+2}} I_{2}^2\bigg( Q^2-C+(2Q(k+2))^{\frac{1}{k+2}}I_{2}\bigg)=0.$

By using $\Gamma_{6}$, the solution is
\begin{equation}
W=\pm\sqrt{Q^2-C+(Q(k+2))^{\frac{2}{k+2}}I_{3}},
\end{equation}
where $I_{3}$ is the constant of integration.

\strut\hfill

{\bf Case:2}

If $k=-2$, then the symmetries of the equation ({\ref{redsecord}}) are
\begin{eqnarray}
\Gamma_{1}&=&\partial_{Q}+\frac{Q}{W}\partial_{W}\\
\Gamma_{2}&=&\frac{(C-Q^{2}+W^{2})}{W}\partial_{W}\\
\Gamma_{3}&=&\frac{Q(C-Q^{2}+W^{2})}{W}\partial_{W}\\
\Gamma_{4}&=& \log[C-Q^2+W^2]\partial_{Q}+\frac{Q\log[C-Q^2+W^2]}{W}\partial_{W}\\
\Gamma_{5}&=&\frac{(C-Q^2+W^2)\log[C-Q^2+W^2]}{W}\partial_{W}\\
\Gamma_{6}&=&4Q\partial_{Q}+\frac{4Q^2+(C-Q^2+W^2)\log[C-Q^2+W^2]}{W}\partial_{W}\\
\Gamma_{7}&=&2Q^2\partial_{Q}+\frac{2Q^3+Q(C-Q^2+W^2)\log[C-Q^2+W^2]}{W}\partial_{W}\\
\Gamma_{8}&=&2Q\log[C-Q^2+W^2]\partial_{Q}\nonumber\\&+&\frac{\log[C-Q^2+W^2](2Q^2+(C-Q^2+W^2)\log[C-Q^2+W^2])}{W}\partial_{W}.
\nonumber\\
\end{eqnarray}

When we use $\Gamma_{1}$, the solution of equation, ({\ref{redsecord}}), is given by
\begin{equation}
W=\pm \sqrt{Q^2+2I_{4}}.
\end{equation}

From $\Gamma_{2}$, the solution is
\begin{equation}
W=\pm \sqrt{(Q^2-C)}.
\end{equation}

From $\Gamma_{3}$, the solution is
\begin{equation}
W=\pm \sqrt{Q^2-C}.
\end{equation}

By using $\Gamma_{4}$ along with the solution obtained from $\Gamma_{1}$, the another solution of $\Gamma_{4}$ is
\begin{equation}
W=\pm\sqrt{1-C+Q^2}.
\end{equation}

By using $\Gamma_{5}$, the solution is
\begin{equation}
W=\pm\sqrt{1-C+Q^2}.
\end{equation}

By using $\Gamma_{6}$, the solution is
\begin{equation}
W=\pm\sqrt{Q^2-C+e^{\pm i{\sqrt{\frac{Q}{I_{5}}}}}}
\end{equation}
with condition $\bigg(e^{ i{\sqrt{\frac{Q}{I_{5}}}}}(C-Q^2)-1\bigg)((k+2)\sqrt{Q}-2i\sqrt{I_{5}})$=$0.$

By using $\Gamma_{7}$, the solution is
\begin{equation}
W=\pm\sqrt{Q^2-C+e^{-\frac{Q}{I_{6}}}}
\end{equation}
with condition $(k+2)\bigg((C-Q^{2})e^{\frac{Q}{I_{6}}}-1\bigg)$=$0.$

By using $\Gamma_{8}$, the solution is
\begin{equation}
W=\pm\sqrt{Q^2-C+e^{-\frac{Q}{I_{7}}}}.
\end{equation}
In the above $I_{4}, I_5, I_6$ and $I_7$ are the constants of integration.

\strut\hfill

In what follows, we reduce the order of another form of the equation (\ref{mch}) by using an equivalence form of the symmetry, $\textbf{X}_{3}$, as follows. Therefore, firstly consider the another form of equation (\ref{mch}),
\begin{eqnarray}
v_{t}+k(t)v^{2}u_{x}+(u^{2}-u_{x}^{2})v_{x}=0,\nonumber\\
v=u-u_{xx}.
\label{Amch}
\end{eqnarray}
 According to equation, $(\ref{Amch})$, the equivalence form of $\textbf{X}_{3}$ is represented as ${\bf\Gamma}_{3}=2t\partial_{t}-u\partial_{u}-v\partial_{v}$. Also the associated transformations for the variables are: $u(x,t)=\frac{y(x)}{\sqrt{t}}$ and $v(x,t)=\frac{z(x)}{\sqrt{t}}$. Therefore the equation (\ref{Amch}) becomes
\begin{eqnarray}
2(y^2-{y^{\prime}}^2)z^{\prime}+2k z^2 y^{\prime}-z=0\nonumber\\
y^{\prime\prime}-y+z=0.\label{SE}
\end{eqnarray}
 The symmetry of the above system $(\ref{SE})$ is $\partial_x$. The reduced ordinary differential system from system $(\ref{SE})$ is given by
 \begin{eqnarray}
2(y^2-p^2)q+2k z^2 p-z=0,\nonumber\\
p p^{\prime}-y+z=0,\label{FE}
\end{eqnarray}
by using the canonical variables $y(x)=y$, $z(x)=z$, $y^{\prime}=p(y)$ and $z^{\prime}=q(z)$ which are obtained from the symmetry $\partial_x$. One can find the solution of (\ref{Amch}) by solving the equation (\ref{FE}).

\section{Painlev\'e Analysis}


Paul Painlev\'e, was French mathematician, was found the Painlev\'e Test. By applying this test one can determine the integrability of a given ordinary differential equations or partial differential equations in terms of analytic functions \cite{Ramani 89, Painleve97, Painleve00, Painleve02}.

\strut\hfill

Mark Feix and his collaborators have investigated the procedure of Ablowitz Ramani Segur (ARS) Algorithm also they found the underlying relation between ARS algorithm and Painlev\'e test \cite{Ablowitz 80a, Ablowitz 80b, Ramani 89}. Based on their outcomes, the solution can be identified as two possible ways(i.e) Right Painleve Series and Left Painleve Series.  The Right painleve series means the Laurent expansion in an ascending form obtained from negative integral power and it constitutes the solution of the given differential equation within a punctured disc centred on the singularity \cite{Andriopoulos 06 a, Krishnakumar 16}. On the other hand, the left painleve series means an expansion descending from the same negative integral power constitutes the solution of the given differential equation shows without the disc \cite{Andriopoulos 06 a, Krishnakumar 16}.

\strut\hfill

The existence of nongeneric negative and  positive resonances was demonstrated by Andriopoulos and Leach in \cite{Andriopoulos 06 a}. They proved that  the solution of the given differential equation in annulus can be studied by standard singularity analysis.

\strut\hfill

In this paper, we can also identify the given differential equation (\ref{mcha}) is integrable, when the ordinary differential equation (\ref{red2}) which obtained from (\ref{mcha}) posses Painlev\'e Property. Also we can find the corresponding solution to the differential equation (\ref{red2}).

 \subsection{Painlev\'e Test}

 Consider the equation (\ref{mcha})
\begin{equation}
u_{t}+k(u-u_{xx})^2u_{x}-u_{xxt}+(u^{2}-{u_{x}}^2)(u_{x}-u_{xxx})=0.
\label{mcha1}
\end{equation}
 As discussed earlier, the above equation (\ref{mcha1}) has been changed to a third-order nonlinear autonomous ordinary differential equation by substituting $r=x-C t$,

\begin{eqnarray}
 Q^{\prime\prime\prime} (C-Q^2+{Q^{\prime}}^2)+k(Q^{\prime\prime}-2Q)Q^{\prime}Q^{\prime\prime}-{Q^{\prime}}^3-(C-Q^2-kQ^2)Q^{\prime}=0, \label{red22}
\end{eqnarray}
where $Q$ is a function of $r$. If we use  this form of (\ref{red22}), it does not pass the Painlev\'e Test. Therefore, we  substitute $Q(r)=\frac{1}{v[r]}$.   Then the equation becomes
 \begin{eqnarray}\label{red21}
 v^2(v^2-C v^4-{v^{\prime}}^2)v^{\prime\prime\prime}+v v^{\prime}v^{\prime\prime}(6C v^4-2(k+3)v^2+2(2k+3){v^{\prime}}^2-k v v^{\prime\prime})&&\nonumber\\-v^{\prime}(2(2k+3){v^{\prime}}^4-C v^6-(4k+7)v^2{v^{\prime}}^2+v^4(1+k+6C {v^{\prime}}^2))=0.
 \end{eqnarray}
 According to the Painlev\'e Test the dominant terms of (\ref{red21}) are given by
\begin{equation}\label{ceq}
v^2(-C v^4-{v^{\prime}}^2)v^{\prime\prime\prime}+v v^{\prime}v^{\prime\prime}(6C v^4+2(2k+3){v^{\prime}}^2-k v v^{\prime\prime})-6C v^4{v^{\prime}}^3-2(2k+3){v^{\prime}}^5
\end{equation}
The behavior of leading-order is $a_{-1} w^{-1}$, where $w=(x-x_0)$.  Next, to find the resonances $(s)$ we substitute $v(w)=a_{-1} w^{-1} + m w^{-1 + s}$ into (\ref{ceq}) and equate the coefficients of $m$ to zero then we have $s=-1, \, 0 ~\text{and} ~1$.

\strut\hfill

 Now we are substituting the expansion $v(w)= - \frac{1}{w} + a_0 + a_1 w $ into (\ref{red21}) to calculate the arbitrariness of $a_0$ and $a_1$. For which collect the coefficients of various  powers of $w$ then equate them to zero. Finally, we get that $a_0~ \text{and} ~a_1$ are all arbitrary constants. Therefore, we conclude that the reduced form of the modified Camassa-Holm Equation (\ref{red21}) passes the Painlev\'{e} Test. As a consequence, the equation (\ref{mcha1}) is obviously an integrable partial differential equation.

\section {Conclusion}

We have examined the Lie point symmetries and Painlev\'e Test for the modified Camassa-Holm Equation with an arbitrary parameter. By the performance of a symmetry analysis we have successfully found the Lie point symmetries of equation (\ref{mcha1}) with the arbitrary parameter $k$. Consequently we have reduced the order of the equation (\ref{mcha1}) from third-order to second-order and then we have also obtained a set of solutions for the resultant second-order ordinary differential equation (\ref{redsecord}). Through the Painlev\'e Test, we were observed that the MCH equation (\ref{mcha1}) is an integrable partial differential equation even with an arbitrary parameter $k$.

\section*{Acknowledgements}
 ADD and KK thank Prof. Stylianos Dimas, S\'ao Jos\'e dos Campos/SP, Brasil, for providing us a new version of the SYM-Package.

\end{document}